You short sold 750 shares of stock at $64 per share at an initial margin of 75%.The maintenance margin is 35%. What is the highest the stock price can go before you receive a margin call? (Express your solution to two decimals of accuracy.)
Answer #1

Initial Margin = 750*64*75% = $36,000

Maintenance Margin = 750*64*35% = $16,800

Hence, loss allowed before margin call = 36000-16800

= $19,200

i.e. Rise in price allowed = 19200/750

= $25.6 per share

Hence, highest price before margin call = 64+25.6

= $89.60 per share